# Responsive Noise-Relaying Diffusion Policy: Responsive and Efficient Visuomotor Control

### v1

@@ Abstract
Imitation learning is an efficient method for teaching robots a variety of tasks. Diffusion Policy, which uses a conditional denoising diffusion process to generate actions, has demonstrated superior performance, particularly in learning from multi-modal demonstrates. However, it relies on executing multiple actions to retain performance and prevent mode bouncing, which limits its responsiveness, as actions are not conditioned on the most recent observations. To address this, we introduce Responsive Noise-Relaying Diffusion Policy (RNR-DP), which maintains a noise-relaying buffer with progressively increasing noise levels and employs a sequential denoising mechanism that generates immediate, noise-free actions at the head of the sequence, while appending noisy actions at the tail. This ensures that actions are responsive and conditioned on the latest observations, while maintaining motion consistency through the noise-relaying buffer. This design enables the handling of tasks requiring responsive control, and accelerates action generation by reusing denoising steps. Experiments on response-sensitive tasks demonstrate that, compared to Diffusion Policy, ours achieves 18% improvement in success rate. Further evaluation on regular tasks demonstrates that RNR-DP also exceeds the best acceleration method by 6.9%, highlighting its computational efficiency advantage in scenarios where responsiveness is less critical.
@@ 1 Introduction
To address these issues, we propose the Responsive Noise-Relaying Diffusion Policy (RNR-DP). This approach fundamentally relies on a noise-relaying buffer that contains increasing noise levels and implements a sequential denoising mechanism. After each denoising step, conditioned on the most recent observations, the model executes an immediate noise-free action at the buffer’s head and appends fully noisy actions at the buffer’s tail. The noise-relaying buffer, combined with the sequential denoising mechanism, not only reuses denoising steps from previous outputs—allowing one denoising step to generate one action—but also ensures active control based on the most recent observations and prevents frequent mode bouncing, maintaining action consistency throughout the entire process. We name it Responsive Noise-Relaying Diffusion Policy because it employs a noise-relaying buffer at its core and has the ability to respond quickly and actively to the environment.
@@ 3 Background
where the initial action sequence ${\mathbf{A}}_{t}$ is sampled from $\mathcal{N}(0,1)$, and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ represents Gaussian noise. The predefined noise schedule functions $\alpha,\beta,$ and $\sigma$ are part of the DDPM scheduler (Ho et al., 2020). Once denoised to ${\mathbf{A}}_{t}^{(0)}$, the agent executes the first $n=T_{a}$ future steps after time $t$.
@@ 4.1 Why Diffusion Policy Needs A Large Action Horizon?
As shown in Chi et al. (2023), Diffusion Policy performs best with a relatively large action horizon $T_{a}$ (e.g., $T_{a}=8$), while $T_{a}=1$ significantly degrades performance. This is because each action sequence is independently denoised from noise, and in multi-modal settings, $T_{a}=1$ allows actions to switch modes, causing inconsistencies. Executing multiple actions within the same mode is crucial for stable performance. To validate this, we train Diffusion Policy on 500 single-modal demonstrations from an RL agent in the ManiSkill2 StackCube task. Results in Table 1 show no performance drop with $T_{a}=1$, even slightly improving over $T_{a}=8$, confirming that mode bouncing in multi-modal settings is the main reason for requiring a large action horizon.
@@ 5 Responsive Noise-Relaying Diffusion Policy
To achieve responsive and efficient control, our proposed approach relies on a noise-relaying buffer and the implemented sequential denoising machanism at its core (Section 5.1) with additional key design choices (Section 5.2). More details of the implementation including policy architecture and hyperparameters are summarized in Appendix B.
@@ 5.1 Noise-Relaying Buffer
The noise-relaying buffer $\mathbf{\tilde{Q}}_{t}=\{\mathbf{a}_{t}^{(1)},\mathbf{a}_{t+1}^{(2)},\dots,\mathbf{a}_{t+f-2}^{(f-1)},\mathbf{a}_{t+f-1}^{(f)}\}$ contains a sequence of noisy actions with linearly increasing noise levels from $1$ to $f$, where $f$ is the buffer capacity as well as the total number of noise levels. As shown in Figure 2, after each denoising step, the trained network transforms $\mathbf{\tilde{Q}}_{t}$ into $\mathbf{Q}_{t}=\{\mathbf{a}_{t}^{(0)},\mathbf{a}_{t+1}^{(1)},\dots,\mathbf{a}_{t+f-2}^{(f-2)},\mathbf{a}_{t+f-1}^{(f-1)}\}$, producing a noise-free action $\mathbf{a}_{t}^{(0)}$ at the head. This clean action is immediately executed, and a fully noisy action is appended to the buffer’s tail. For the next step, the buffer reuses $f-1$ denoising steps from previous outputs, ensuring consistency and avoiding full denoising from scratch. This sequential denoising mechanism conditions clean actions on the latest observations, enabling responsive and long-term active control. Pseudocode is provided in Algorithm 1.
Figure 2: Inference Overview of Responsive Noise-Relaying Diffusion Policy (RNR-DP). The core of RNR-DP is the noise-relaying buffer (Section 5.1) and it has 3 stages during the entire control-loop, as exemplified by the transition between time step $t$ and time step $t+1$, (1) The buffer contains noisy actions with increasing noise levels (2) After denoising once, each action in the buffer is denoised for one step, clean action at the buffer’s head is removed and executed (dequeue) (3) The remaining noisy actions are left shifted for one slot and a fully noisy action is appended to the buffer’s tail (enqueue). The conditioning data is discussed with more details in Section 5.2 and Section B.3.
@@ 5.2 Key Design Choices
Mixture Noise Scheduling We train the denoiser network following the DDPM (Ho et al., 2020) framework, and allow each action in $\mathbf{A}_{t}$ perturbed by independent noise levels. Given a fixed variance schedule $\beta_{1},\ldots,\beta_{f}$ ($\beta_{1}<\beta_{f}$), any $\mathbf{a}_{j}$ in $\mathbf{A}_{t}$ can be perturbed by one of the $f$ levels. During training, we use a mixed per-action noise injection scheme (mixture schedule): with probability $p_{\mathrm{linear}}$, actions are perturbed by linearly increasing variances (linear schedule); with probability $1-p_{\mathrm{linear}}$, actions are perturbed by random variances from $\beta_{1}$ to $\beta_{f}$ (random schedule). The random schedule trains the model to denoise actions independently $p_{\theta}(\mathbf{a}^{(k-1)}\mid\mathbf{a}^{(k)};\mathbf{O}),1\leq k\leq f$, while the linear schedule ensures smooth transitions across consecutive actions during inference. Unlike Diffusion Policy, which applies a single variance level to all actions in $\mathbf{A}_{t}$ per iteration, our mixture schedule enables diverse and robust training.
@@
Laddering Initialization During inference, we use a noise-relaying buffer with $f$ noisy action frames, following Algorithm 1. Initially, the buffer contains $f$ fully noisy actions sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e., $\tilde{\mathbf{Q}}=\{\mathbf{z}_{j}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\mid j=1,\ldots,f\}$. To align with training and avoid performance drops, we initialize the buffer by iteratively denoising it $f$ times using the random schedule conditioned on the initial observation $\mathbf{O}_{0}$. This transforms the buffer to follow the linear schedule, ensuring smooth and responsive control. Since this process transitions the buffer from uniform noise to monotonically increasing variances like a ladder, we term it laddering initialization, as illustrated in Figure 3.
Noise-Aware Conditioning Unlike Diffusion Policy (Chi et al., 2023), which encodes a single diffusion step to one time embedding, our approach uses mixed scheduling and a noise-relaying buffer to handle multiple noise levels, requiring awareness of multiple diffusion steps. For each action’s noise level $k_{j}\ (1\leq k_{j}\leq f)$, we use an MLP to encode a time embedding, yielding $f$ embeddings. These are appended to the observation features from the encoder $E_{\mathrm{obs}}$. This allows the buffer, during both training and inference, to decode actions at each noise level using up-to-date observations, enabling more dynamic and consistent behaviors. See Section B.3 for more details.
@@ 6 Experiments
    Can Responsive Noise-Relaying Diffusion Policy outperform Diffusion Policy by delivering more responsive control? (Section 6.3)?
@@
    Can Responsive Noise-Relaying Diffusion Policy function as a superior acceleration method compared to commonly used alternatives on simpler tasks that do not require responsive control (Section 6.4)?
@@
    What are the effects of the components introduced by Responsive Noise-Relaying Diffusion Policy (Section 6.5)?
@@ 6.1 Experimental Setup
Figure 4: Task Visualization in Simulation. ManiSkill tasks including PushT, RollBall, PushChair, StackCube, TurnFaucet; Adroit tasks including Relocate, Door, Pen, Hammer. See Section A.1 for more properties about each task.
@@ 6.1.1 Task Descriptions
Our experiments are conducted on 9 tasks from 3 benchmarks: ManiSkill2 (robotic manipulation; 3 tasks), ManiSkill3 (robotic manipulation; 2 tasks) and Adroit (dexterous manipulation; 4 tasks). These tasks are separated into 2 groups to validate the responsiveness and efficiency of Responsive Noise-Relaying Diffusion Policy.
Response-sensitive Group: Tasks Involving Dynamic Object Manipulation Requiring Responsive Control We consider 5 challenging tasks involving contact-rich, dynamic object manipulation to assess the responsiveness of RNR-DP. PushT requires using a stick to push a T-shaped block to a target location and orientation. RollBall involves pushing and rolling a ball to a randomized goal region. PushChair tests bimanual manipulation of articulated objects with variations. Adroit Relocate involves picking up a ball and moving it to a goal position, while Adroit Door requires unlocking and opening a door. See Figure 4 for task visualizations, and more details are included in Section A.1.
Regular Group: Simpler Tasks that Do Not Require Responsive Control We consider the rest 4 simpler tasks that do not require responsive control to validate the efficiency of RNR-DP. StackCube requires picking up a cube and stack it onto another cube. TurnFaucet uses a stationary arm to turn on faucets of various geometries and topology. Adroit Pen repositions the blue pen to match the orientation of the green target. Adroit Hammer picks up a hammer and drives a nail into a board. More details are included in Section A.1.
@@ 6.2 Baselines
Streaming Diffusion Policy (SDP) is a recent advancement over Diffusion Policy that stays close to our approach. See Section C.2 for detailed discussion.
@@ 6.3 Results & Analysis on Responsive Control
First, we provide comprehensive evaluation and validate the responsiveness of RNR-DP on 5 tasks (Response-sensitive Group) involving contact-rich dynamic object manipulation, a primary focus of our policy. Figure 1 shows an overview of performance improvement. As shown in Table 2 and Table 3, RNR-DP consistently outperforms Diffusion Policy across all tested scenarios in both state and visual experiments. Specifically, in state experiments, RNR-DP achieves a $15.1\%$ improvement over Diffusion Policy, while in visual experiments, it demonstrates a $24.9\%$ improvement. Overall, RNR-DP surpasses Diffusion Policy by $18.0\%$. Notably, in the Relocate task, RNR-DP achieves a significant performance boost of $38.6\%$ over Diffusion Policy.
@@
Table 3: Evaluation on tasks (Response-sensitive Group) from ManiSkill and Adroit benchmarks (Visual Observations). We report values under the same settings as in Table 2. Tasks in which none of the methods achieve a reasonable success rate under visual observations are omitted.
@@ 6.4 Results & Analysis on Efficient Control
In this section, we conduct extended experiments on 4 simpler tasks (Regular Group) that do not require responsive control to evaluate the efficiency of RNR-DP. An overview of the performance improvements is provided in Figure 1. Section 6.4.1 introduces a metric designed for fair efficiency comparisons, while Section 6.4.2 provides a detailed analysis of the empirical results compared to commonly used acceleration methods.
@@ 6.4.1 Neural Function Evaluations per Action (NFEs/a)
Adopting the settings from Chi et al. (2023), Diffusion Policy and related methods utilize a relatively large action horizon ($T_{a}$) to achieve better performance, whereas RNR-DP employs a single-action rollout ($T_{a}=1$) at each inference. To facilitate a fair comparison of efficiency, we introduce a new metric, Neural Function Evaluations per Action (NFEs/a), as defined in Equation 1.
@@ 6.4.2 Empirical Comparison with Commonly Used Acceleration Methods
To evaluate the efficiency of Responsive Noise-Relaying Diffusion Policy (RNR-DP), we compare it against common acceleration methods, including DDIM (Chi et al., 2023) and Consistency Policy (Prasad et al., 2024), on simpler tasks that do not require responsive control (Regular Group). As shown in Table 4 and Section C.1, RNR-DP consistently achieves the highest overall performance among all DDIM and Consistency Policy variants. For a fair comparison, we specifically evaluate RNR-DP against 8-step DDIM (NFEs/a = 1) and 8-step-chaining Consistency Policy (NFEs/a = 1). In state-based experiments, RNR-DP outperforms 8-step DDIM by 5.7% and 8-step-chaining Consistency Policy by 66.2%. In vision-based experiments, it surpasses 8-step DDIM by 8.5% and 8-step-chaining Consistency Policy by 3.4%, yielding an overall improvement of 6.9% over 8-step DDIM and 32.0% over 8-step-chaining Consistency Policy. Additionally, both DDIM and Consistency Policy struggle in certain tasks. For instance, 8-step DDIM achieves a $0\%$ success rate on Adroit Hammer, while Consistency Policy performs poorly on StackCube and TurnFaucet. In contrast, RNR-DP demonstrates consistent and robust performance across all tasks. Notably, RNR-DP achieves an average success rate comparable to Diffusion Policy across all state and visual experiments while being 12.5 times faster. These results confirm that even on simpler tasks, RNR-DP serves as a superior acceleration method, enabling efficient control.

### v2

@@ Abstract
Imitation learning is an efficient method for teaching robots a variety of tasks. Diffusion Policy, which uses a conditional denoising diffusion process to generate actions, has demonstrated superior performance, particularly in learning from multi-modal demonstrates. However, it relies on executing multiple actions predicted from the same inference step to retain performance and prevent mode bouncing, which limits its responsiveness, as actions are not conditioned on the most recent observations. To address this, we introduce Responsive Noise-Relaying Diffusion Policy (RNR-DP), which maintains a noise-relaying buffer with progressively increasing noise levels and employs a sequential denoising mechanism that generates immediate, noise-free actions at the head of the sequence, while appending noisy actions at the tail. This ensures that actions are responsive and conditioned on the latest observations, while maintaining motion consistency through the noise-relaying buffer. This design enables the handling of tasks requiring responsive control, and accelerates action generation by reusing denoising steps. Experiments on response-sensitive tasks demonstrate that, compared to Diffusion Policy, ours achieves 18% improvement in success rate. Further evaluation on regular tasks demonstrates that RNR-DP also exceeds the best acceleration method (DDIM) by 6.9% in success rate, highlighting its computational efficiency advantage in scenarios where responsiveness is less critical. Our project page is available at https://rnr-dp.github.io/.
@@ 1 Introduction
To address these issues, we propose the Responsive Noise-Relaying Diffusion Policy (RNR-DP). We define responsiveness as the degree to which the current action leverages the most recent observation. Our approach fundamentally relies on a noise-relaying buffer that contains increasing noise levels and implements a sequential denoising mechanism. After each denoising step, conditioned on the most recent observations, the model executes an immediate noise-free action at the buffer’s head and appends fully noisy actions at the buffer’s tail. The noise-relaying buffer, combined with the sequential denoising mechanism, not only reuses denoising steps from previous outputs—allowing one denoising step to generate one action—but also ensures active control based on the most recent observations and prevents frequent mode bouncing, maintaining action consistency throughout the entire process. We name it Responsive Noise-Relaying Diffusion Policy because it employs a noise-relaying buffer at its core and has the ability to respond quickly and actively to the environment.
@@ 3 Background
where the initial action sequence ${\mathbf{A}}_{t}$ is sampled from $\mathcal{N}(0,1)$, and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ represents Gaussian noise. The predefined noise schedule functions $\alpha,\gamma,$ and $\sigma$ are part of the DDPM scheduler (Ho et al., 2020). Once denoised to ${\mathbf{A}}_{t}^{(0)}$, the agent executes the first $n=T_{a}$ future steps after time $t$.
@@ 4.1 Why Diffusion Policy Needs A Large Action Horizon?
As shown in Chi et al. (2023), Diffusion Policy performs best with a relatively large action horizon $T_{a}$ (e.g., $T_{a}=8$), while $T_{a}=1$ significantly degrades performance. This is because each action sequence is independently denoised from noise, and in multi-modal settings, $T_{a}=1$ allows actions to switch modes, causing inconsistencies. Executing multiple actions within the same mode is crucial for stable performance. To validate this, we train Diffusion Policy on 500 single-modal demonstrations from an RL agent in the ManiSkill2 StackCube task. Results in Table˜1 show no performance drop with $T_{a}=1$, even slightly improving over $T_{a}=8$, confirming that mode bouncing in multi-modal settings is the main reason for requiring a large action horizon.
@@ 5 Responsive Noise-Relaying Diffusion Policy
To achieve responsive and efficient control, our proposed approach relies on a noise-relaying buffer and the implemented sequential denoising machanism at its core (Section˜5.1) with additional key design choices (Section˜5.2). More details of the implementation including policy architecture and hyperparameters are summarized in Appendix˜B.
@@ 5.1 Noise-Relaying Buffer
The noise-relaying buffer $\mathbf{\tilde{Q}}_{t}=\{\mathbf{a}_{t}^{(1)},\mathbf{a}_{t+1}^{(2)},\dots,\mathbf{a}_{t+f-2}^{(f-1)},\mathbf{a}_{t+f-1}^{(f)}\}$ contains a sequence of noisy actions with linearly increasing noise levels from $1$ to $f$, where $f$ is the buffer capacity as well as the total number of noise levels. As shown in Figure˜2, after each denoising step, the trained network transforms $\mathbf{\tilde{Q}}_{t}$ into $\mathbf{Q}_{t}=\{\mathbf{a}_{t}^{(0)},\mathbf{a}_{t+1}^{(1)},\dots,\mathbf{a}_{t+f-2}^{(f-2)},\mathbf{a}_{t+f-1}^{(f-1)}\}$, producing a noise-free action $\mathbf{a}_{t}^{(0)}$ at the head. This clean action is immediately executed, and a fully noisy action is appended to the buffer’s tail. For the next step, the buffer reuses f1f-1 denoising steps from previous outputs, ensuring consistency and avoiding full denoising from scratch. This sequential denoising mechanism conditions clean actions on the latest observations, enabling responsive and long-term active control. Pseudocode is provided in Algorithm˜1.
Figure 2: Inference Overview of Responsive Noise-Relaying Diffusion Policy (RNR-DP). The core of RNR-DP is the noise-relaying buffer (Section˜5.1) and it has 3 stages during the entire control-loop, as exemplified by the transition between time step $t$ and time step $t+1$, (1) The buffer contains noisy actions with increasing noise levels (2) After denoising once, each action in the buffer is denoised for one step, clean action at the buffer’s head is removed and executed (dequeue) (3) The remaining noisy actions are left shifted for one slot and a fully noisy action is appended to the buffer’s tail (enqueue). The conditioning data is discussed with more details in Section˜5.2 and Section˜B.3.
@@ 5.2 Key Design Choices
Mixture Noise Scheduling We train the denoiser network following the DDPM (Ho et al., 2020) framework, and allow each action in $\mathbf{A}_{t}$ perturbed by independent noise levels. Given a fixed variance schedule $\beta_{1},\ldots,\beta_{f}$ ($\beta_{1}<\beta_{f}$), any $\mathbf{a}_{j}$ in $\mathbf{A}_{t}$ can be perturbed by one of the $f$ levels. During training, we use a mixed per-action noise injection scheme (mixture schedule): with probability $p_{\mathrm{linear}}$, actions are perturbed by linearly increasing variances (linear schedule); with probability $1-p_{\mathrm{linear}}$, actions are perturbed by random variances from $\beta_{1}$ to $\beta_{f}$ (random schedule). The random schedule trains the model to denoise actions independently $p_{\theta}(\mathbf{a}^{(k-1)}\mid\mathbf{a}^{(k)};\mathbf{O}),1\leq k\leq f$, while the linear schedule ensures smooth transitions across consecutive actions during inference. Unlike Diffusion Policy, which applies a single variance level to all actions in $\mathbf{A}_{t}$ per iteration, our mixture schedule enables diverse and robust training. An illustrative visualization is included in Appendix˜I.
@@
Laddering Initialization During inference, we use a noise-relaying buffer with $f$ noisy action frames, following Algorithm˜1. Initially, the buffer contains $f$ fully noisy actions sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e., $\tilde{\mathbf{Q}}=\{\mathbf{z}_{j}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\mid j=1,\ldots,f\}$. To align with training and avoid performance drops, we initialize the buffer by iteratively denoising it $f$ times using the random schedule conditioned on the initial observation $\mathbf{O}_{0}$. This transforms the buffer to follow the linear schedule, ensuring smooth and responsive control. Since this process transitions the buffer from uniform noise to monotonically increasing variances like a ladder, we term it laddering initialization, as illustrated in Figure˜3. A more detailed illustrative visualization is included in Appendix˜J.
Noise-Aware Conditioning Unlike Diffusion Policy (Chi et al., 2023), which encodes a single diffusion step to one time embedding, our approach uses mixed scheduling and a noise-relaying buffer to handle multiple noise levels, requiring awareness of multiple diffusion steps. For each action’s noise level $k_{j}\ (1\leq k_{j}\leq f)$, we use an MLP to encode a time embedding, yielding $f$ embeddings. These are appended to the observation features from the encoder $E_{\mathrm{obs}}$. This allows the buffer, during both training and inference, to decode actions at each noise level using up-to-date observations, enabling more dynamic and consistent behaviors. See Section˜B.3 for more details.
@@ 6 Experiments
    Can Responsive Noise-Relaying Diffusion Policy outperform Diffusion Policy by delivering more responsive control? (Section˜6.3)?
@@
    Can Responsive Noise-Relaying Diffusion Policy function as a superior acceleration method compared to commonly used alternatives on simpler tasks that do not require responsive control (Section˜6.4)?
@@
    What are the effects of the components introduced by Responsive Noise-Relaying Diffusion Policy (Section˜6.5)?
@@ 6.1 Experimental Setup
Figure 4: Task Visualization in Simulation. ManiSkill tasks including PushT, RollBall, PushChair, StackCube, TurnFaucet; Adroit tasks including Relocate, Door, Pen, Hammer. See Section˜A.1 for more properties about each task.
@@ 6.1.1 Task Descriptions
Our experiments are conducted on 9 tasks from 3 benchmarks: ManiSkill2 (robotic manipulation; 3 tasks), ManiSkill3 (robotic manipulation; 2 tasks) and Adroit (dexterous manipulation; 4 tasks). These tasks are separated into 2 groups to validate the responsiveness and efficiency of Responsive Noise-Relaying Diffusion Policy. A detailed discussion of task groupings and classification criteria can be found in Appendix˜D.
Response-sensitive Group: Tasks Involving Dynamic Object Manipulation Requiring Responsive Control We consider 5 challenging tasks involving contact-rich, dynamic object manipulation to assess the responsiveness of RNR-DP. PushT requires using a stick to push a T-shaped block to a target location and orientation. RollBall involves pushing and rolling a ball to a randomized goal region. PushChair tests bimanual manipulation of articulated objects with variations. Adroit Relocate involves picking up a ball and moving it to a goal position, while Adroit Door requires unlocking and opening a door. See Figure˜4 for task visualizations, and more details are included in Section˜A.1.
Regular Group: Simpler Tasks that Do Not Require Responsive Control We consider the rest 4 simpler tasks that do not require responsive control to validate the efficiency of RNR-DP. StackCube requires picking up a cube and stack it onto another cube. TurnFaucet uses a stationary arm to turn on faucets of various geometries and topology. Adroit Pen repositions the blue pen to match the orientation of the green target. Adroit Hammer picks up a hammer and drives a nail into a board. More details are included in Section˜A.1.
@@ 6.2 Baselines
Streaming Diffusion Policy (SDP) is a recent advancement over Diffusion Policy that stays close to our approach. See Section˜C.2 for detailed discussion.
@@ 6.3 Results & Analysis on Responsive Control
First, we provide comprehensive evaluation and validate the responsiveness of RNR-DP on 5 tasks (Response-sensitive Group) involving contact-rich dynamic object manipulation, a primary focus of our policy. Figure˜1 shows an overview of performance improvement. As shown in Table˜2 and Table˜3, RNR-DP consistently outperforms Diffusion Policy across all tested scenarios in both state and visual experiments. Specifically, in state experiments, RNR-DP achieves a $15.1\%$ improvement over Diffusion Policy, while in visual experiments, it demonstrates a $24.9\%$ improvement. Overall, RNR-DP surpasses Diffusion Policy by $18.0\%$. Notably, in the Relocate task, RNR-DP achieves a significant performance boost of $38.6\%$ over Diffusion Policy.
@@
Table 3: Evaluation on tasks (Response-sensitive Group) from ManiSkill and Adroit benchmarks (Visual Observations). We report values under the same settings as in Table˜2. Tasks in which none of the methods achieve a reasonable success rate under visual observations are omitted.
@@ 6.4 Results & Analysis on Efficient Control
In this section, we conduct extended experiments on 4 simpler tasks (Regular Group) that do not require responsive control to evaluate the efficiency of RNR-DP. An overview of the performance improvements is provided in Figure˜1. Section˜6.4.1 introduces a metric designed for fair efficiency comparisons, while Section˜6.4.2 provides a detailed analysis of the empirical results compared to commonly used acceleration methods.
@@ 6.4.1 Neural Function Evaluations per Action (NFEs/a)
Adopting the settings from Chi et al. (2023), Diffusion Policy and related methods utilize a relatively large action horizon ($T_{a}$) to achieve better performance, whereas RNR-DP employs a single-action rollout ($T_{a}=1$) at each inference. To facilitate a fair comparison of efficiency, we introduce a new metric, Neural Function Evaluations per Action (NFEs/a), as defined in Equation˜1.
@@ 6.4.2 Empirical Comparison with Commonly Used Acceleration Methods
To evaluate the efficiency of Responsive Noise-Relaying Diffusion Policy (RNR-DP), we compare it against common acceleration methods, including DDIM (Chi et al., 2023) and Consistency Policy (Prasad et al., 2024), on simpler tasks that do not require responsive control (Regular Group). As shown in Table˜4 and Section˜C.1, RNR-DP consistently achieves the highest overall performance among all DDIM and Consistency Policy variants. For a fair comparison, we specifically evaluate RNR-DP against 8-step DDIM (NFEs/a = 1) and 8-step-chaining Consistency Policy (NFEs/a = 1). In state-based experiments, RNR-DP outperforms 8-step DDIM by 5.7% and 8-step-chaining Consistency Policy by 66.2%. In vision-based experiments, it surpasses 8-step DDIM by 8.5% and 8-step-chaining Consistency Policy by 3.4%, yielding an overall improvement of 6.9% over 8-step DDIM and 32.0% over 8-step-chaining Consistency Policy. Additionally, both DDIM and Consistency Policy struggle in certain tasks. For instance, 8-step DDIM achieves a $0\%$ success rate on Adroit Hammer, while Consistency Policy performs poorly on StackCube and TurnFaucet. In contrast, RNR-DP demonstrates consistent and robust performance across all tasks. Notably, RNR-DP achieves an average success rate comparable to Diffusion Policy across all state and visual experiments while being 12.5 times faster. These results confirm that even on simpler tasks, RNR-DP serves as a superior acceleration method, enabling efficient control.